OpenStudy (anonymous):

(sqrt 3 - sqrt 2)(5 Sqrt 2 - 4 sqrt 3)

OpenStudy (anonymous):

This question isn't hard. just FOIL it's going to be messy, so just be careful with each step (a-b)(c-d) ac - bc - ad + bd good luck!

OpenStudy (radar):

\[(\sqrt{3}-\sqrt{2})(5\sqrt{2}-4\sqrt{3})=5\sqrt{6}-10-12+4\sqrt{6}=9\sqrt{6}-22\]

OpenStudy (anonymous):

lols go radar!

OpenStudy (radar):

Review the rules for multiplication of radicals\[\sqrt{ab}=\sqrt{a}\sqrt{b}\]

OpenStudy (radar):

\[3\sqrt{2} \times4\sqrt{6}=12\sqrt{12}\]

OpenStudy (radar):

If terms inside radicals are identical, the radicals can be added or subtracted as required.

OpenStudy (anonymous):

thanks